# Exploring Temporally-Aware Features for Point Tracking

### v1

@@ Abstract
Figure 1: Chrono is a highly precise, temporally-aware feature backbone specifically designed for point tracking. Without relying on extensive iterative refinement [19, 12, 13, 11, 25] or test-time optimization [45, 47], our model demonstrates a significant improvement over existing feature backbones such as TSM-ResNet [31].
@@ 1 Introduction
Despite its critical role, temporally-aware backbone has been relatively under-explored in this area. Early efforts, such as simply incorporating TSM-ResNet [31] in point tracking models [12, 13], yielded limited improvements. This outcome may be attributed to two factors: First, training on small synthetic datasets [16] may limit generalization to real-world data. Second, TSM-ResNet’s temporal awareness, constrained to adjacent frames, may not provide long-term temporal context for point tracking.
@@
To address these limitations, we propose a feature backbone for point tracking that incorporates temporal awareness, dubbed as Chrono. Our key insights are twofold: First, we leverage robust feature representations learned from large-scale, real-world datasets. We employ DINOv2 [36], a pre-trained model known for strong feature representations across various tasks such as segmentation, classification, and localization [29, 36]. However, these representations are not directly compatible with point tracking due to a lack of temporal awareness. Second, we incorporate temporal awareness by designing a temporal adapter that enables the pre-trained backbone to process data with temporal context, without losing its learned knowledge. This enables the use of these strong feature backbones for point tracking tasks. Unlike the adjacent-frame context used in TSM-ResNet [31], our approach incorporates a temporal context that is six times longer, enabling the capture of complex dynamics and enhancing tracking performance.
Experimental results show that tracks estimated from Chrono features, computed solely using soft argmax without any learnable layers or additional temporal information after feature extraction, significantly outperform traditional feature backbones commonly used in point tracking as well as DINOv2 [36]. Specifically, Chrono achieves a +20.6%p increase in the position accuracy on the TAP-Vid-DAVIS [12] dataset compared to TSM-ResNet-18 [31]. Additionally, our method demonstrates superior efficiency and comparable performance compared to point tracking models equipped with refiners that inject temporal information post-feature extraction. Despite the absence of learnable layers after feature extraction, Chrono achieves high precision, underscoring that directly embedding temporal information within the features is both efficient and powerful. This finding demonstrates that Chrono models temporal dynamics and motion as effectively as more complex refiner-based methods.
@@ 2 Related Work
With recent advances in large-scale self-supervised training [22, 36], effectively transferring its vast knowledge has become a key challenge. Specifically, the strong features of DINOv2 have been adapted to various tasks. [42] fine-tunes DINO for video tasks in a self-supervised manner. Additionally, DINO has been applied to various correspondence tasks [1, 34, 43, 29, 17] and has demonstrated robustness in establishing correspondence and semantic segmentation [18, 1, 35], showcasing its rich semantics. We focus on adapting this representation to the point tracking task, enhancing its temporal consistency while preserving its pre-trained knowledge.
@@ 3.1 Preliminaries and Motivation
Point tracking [30, 13, 14, 12, 25, 11, 10] in videos involves establishing correspondences of specific points across frames, determining their positions over time, and identifying whether they are visible or occluded. Formally, given a video sequence $\{\mathcal{I}_{t}\}_{t=0}^{T-1}$, where $\mathcal{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ represents the $t$-th frame of height $H$ and width $W$, and a query point $q=(x_{q},y_{q},t_{q})\in\mathbb{R}^{3}$ at frame $t_{q}$, the goal is to produce a trajectory $\mathcal{T}=\{\hat{p}_{t}\}_{t=0}^{T-1}$, where $\hat{p}_{t}\in\mathbb{R}^{2}$ denotes the estimated position of the point at time $t$, and an associated visibility probability $\mathcal{V}=\{v_{t}\}_{t=0}^{T-1}$, where $v_{t}\in[0,1]$ indicates the likelihood of the point being visible at frame $t$.
Motivation and overview.
@@ 3.2 Temporal Adapter
$\alpha^{(t,n)}(x,y)=\frac{\exp\left(\mathbf{Q}_{t}(x,y)\cdot\mathbf{K}_{t+n}(x,y)\right)}{\sum_{n^{\prime}=-k}^{k}\exp\left(\mathbf{Q}_{t}(x,y)\cdot\mathbf{K}_{t+n^{\prime}}(x,y)\right)},$ (3)
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are projections of $\mathbf{f}^{\text{down}}$, obtained through the linear projection layers $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$.
@@ 3.3 Point Prediction with the Feature Backbone
To estimate the position of points in frame $t$, we apply the soft-argmax operation to the correlation map $C_{t}$. The soft-argmax computes a weighted average of all spatial positions, with weights given by the softmax of the correlation values:
@@
$\mathcal{L}_{\text{Huber}}(\hat{\mathbf{p}}_{t},\mathbf{p}_{t})=\begin{cases}\frac{1}{2}\left\|\hat{\mathbf{p}}_{t}-\mathbf{p}_{t}\right\|^{2},\!&\text{if }\left\|\hat{\mathbf{p}}_{t}-\mathbf{p}_{t}\right\|\leq\delta,\\ \delta\cdot\left(\left\|\hat{\mathbf{p}}_{t}-\mathbf{p}_{t}\right\|-\frac{1}{2}\delta\right),&\text{otherwise},\end{cases}$ (9)
where $\mathbf{p}_{t}$ is the ground truth position, and $\delta$ is a threshold parameter. For occluded points indicated by the ground truth visibility status $v_{t}$, we exclude the loss computation, effectively masking out those time steps:
$\mathcal{L}_{t}=(1-v_{t})\cdot\mathcal{L}_{\text{Huber}}(\hat{\mathbf{p}}_{t},\mathbf{p}_{t}).$ (10)
@@ 4 Experiments
Backbone Kinetics-Strided DAVIS-Strided
$<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$
ResNet-18 [20, 13] 10.5 35.3 65.7 81.3 88.6 56.3 9.7 31.1 60.6 78.3 87.0 53.3
TSM-ResNet-18 [31, 12] 9.15 33.2 64.6 79.2 86.5 54.5 8.2 26.8 53.6 73.5 83.9 49.2
DINOv2 (ViT-S/14) [36] 4.9 15.8 41.9 73.8 86.0 44.5 7.3 22.7 52.9 80.0 89.5 50.4
DINOv2 (ViT-B/14) [36] 5.9 18.6 45.9 75.7 86.9 46.6 10.0 28.6 59.5 82.9 87.4 54.4
Chrono (ViT-S/14) 32.2 55.2 73.9 84.0 88.7 66.8 29.7 56.4 76.7 86.8 90.9 68.0
Chrono (ViT-B/14) 33.5 57.2 75.8 85.8 90.2 68.5 31.8 59.2 78.8 88.4 92.2 70.1
Table 1: Quantitative comparison on the TAP-Vid datasets [12] with the strided query mode. Compared to DINOv2 and previous feature backbones used in point tracking, our model shows a significant performance boost.
Backbone Kinetics-First DAVIS-First
$<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$
ResNet-18 [20, 13] 8.6 28.8 56.5 74.2 83.3 50.3 9.0 27.3 54.9 73.7 84.1 49.8
TSM-ResNet-18 [31, 12] 7.8 28.1 55.2 71.4 80.2 48.6 7.3 23.1 46.7 66.6 79.2 44.6
DINOv2 (ViT-S/14) [36] 4.2 13.6 36.3 65.9 79.2 39.8 6.0 19.9 47.5 73.7 84.5 46.3
DINOv2 (ViT-B/14) [36] 5.1 16.0 40.0 67.9 80.4 41.9 8.9 24.7 53.8 77.0 87.1 50.3
Chrono (ViT-S/14) 24.8 46.2 65.8 77.5 83.1 59.5 24.0 49.2 71.2 82.8 87.9 63.0
Chrono (ViT-B/14) 26.0 48.4 68.2 79.8 85.3 61.6 26.1 52.6 74.5 84.9 90.0 65.6
Table 2: Quantitative comparison on the TAP-Vid datasets [12] with the query first mode. Compared to other feature backbones widely used in point tracking, our model shows exceptional performance in position accuracy.
@@ 4.2 Evaluation Protocol
To assess our approach, we employ the TAP-Vid benchmark [12], which comprises both real and synthetic video sets. The real videos come with accurate annotation tracks, while synthetic videos are paired with perfect ground-truth trajectories. For our analysis, we focus on the real video datasets, specifically TAP-Vid-DAVIS and TAP-Vid-Kinetics. The DAVIS dataset [38] provides 30 videos that include challenges such as substantial scale shifts of objects, while Kinetics [27] consists of 1,189 YouTube videos featuring diverse difficulties, including intense motion blur and abrupt scene transitions.
@@
We assess the accuracy of the predicted tracks using two metrics: position accuracy at various thresholds and average position accuracy $(<\delta^{x}_{avg})$. Position accuracy is evaluated at five threshold levels: $<\delta^{0}$, $<\delta^{1}$, $<\delta^{2}$, $<\delta^{3}$, and $<\delta^{4}$, representing accuracies within pixel distances of 1, 2, 4, 8, and 16, respectively. Each $<\delta^{x}$ score indicates the percentage of visible ground-truth points with predictions falling within the specified threshold. $<\delta^{x}_{avg}$ averages these scores to provide an overall measure of position accuracy across all thresholds.
@@ 4.3 Main Results
We evaluate our method against the backbones [20, 31] commonly used in point tracking [11, 12, 13, 25, 26] and DINOv2 [36] in both the strided query (Table 1) and the first query modes (Table 2). Specifically, we compare with ResNet-18 [20] from TAPIR [13] and TSM-ResNet-18 [31] from TAP-Net [12], both trained for point tracking. We report position accuracy to demonstrate the effectiveness of temporal information in our model.
Our small model, Chrono (ViT-S/14) with DINOv2 (ViT-S/14), achieves state-of-the-art performance in position accuracy across all thresholds, surpassing other backbones. Chrono (ViT-B/14) with DINOv2 (ViT-B/14) delivers even better results. Across both Kinetics and DAVIS datasets, in the strided and first query modes, Chrono (ViT-S/14) and Chrono (ViT-B/14) consistently outperform ResNet-18, TSM-ResNet-18, and other DINOv2 variants in PCK accuracy at all thresholds. The high accuracy at $<\delta^{0}$ underscores the precision of Chrono features in point tracking, with further gains at higher thresholds. At $<\delta^{4}$, only Chrono backbones achieve over 90% accuracy on the DAVIS dataset in strided mode. Both Chrono (ViT-S/14) and Chrono (ViT-B/14) reach the highest position accuracy among backbones, demonstrating Chrono ’s ability to effectively model spatial and temporal information in videos.
@@ 4.4 Analysis and Ablation Study
Comparison to point tracking with iterative refiner.
As shown in Table 3, Chrono, although solely a feature backbone without a refiner, achieves point estimation precision comparable to full pipelines with refiners, while demonstrating significantly higher efficiency.
Unlike other models, which require refiners to inject temporal information for each query point, Chrono embeds temporal awareness directly in the feature backbone. This approach significantly increases throughput, with Chrono (ViT-B/14) achieving 12.5 times the speed of TAPIR while maintaining only a 3.5%p difference in precision, and Chrono (ViT-S/14) delivering 16.4 times the throughput. These results demonstrate that Chrono provides a simple yet highly effective solution, offering performance comparable to refiner-based models without requiring iterative processing. We define the number of learnable parameters after feature extraction as the number of refiner parameters.
Figure 5: Visualization of features. We visualize the features from Chrono and those from DINOv2 with PCA. Our model demonstrates better temporal smoothness than DINOv2.
Method $<\delta^{x}_{avg}$ Throughput # of Refiner
(points/sec) Params
RAFT [44] 46.3 23,405.71 4.2M
TAP-Net [12] 53.1 29,535.98 5.5K
TAPIR [13] 73.6 2,097.32 25.9M
PIPs [19] 59.4 46.43 26.0M
Chrono (ViT-S/14) 68.0 34,396.30 None
Chrono (ViT-B/14) 70.1 26,139.86 None
Table 3: Comparison on point tracking pipelines using refiners. Although Chrono does not use a heavy iterative refiner, it achieves comparable performance with significantly higher throughput. All metrics are measured on a 24-frame video using a single NVIDIA RTX 3090 GPU, and the TAP-Vid-DAVIS dataset [12].
@@
Analysis on Chrono as a backbone in existing point tracking pipeline.
@@
To model temporal information in our temporal adapter, we tested two approaches for temporal aggregation: convolution and attention. In Table 5, we present an ablation study where only the temporal aggregation layer within Chrono is modified, comparing 1D convolution, 3D convolution, and 1D attention, with all other components kept constant. While 1D convolution and 1D attention aggregate along the temporal axis only, 3D convolution considers both spatial and temporal dimensions. Due to computational constraints, this study was conducted on Chrono (ViT-S/14).
@@
Early transformer blocks often capture local details, while later blocks tend to focus on broader, global patterns and complex relationships [15, 39]. Therefore, it’s plausible that adapters in the initial blocks may learn local motion cues, while those in later blocks might capture overall motion patterns. Placing adapters only in the early or later blocks might emphasize either local or global motion, respectively. While the alternating configuration shows some improvement, it may not be optimal. Integrating temporal adapters between each block could potentially allow for a more balanced capture of both local and global motion patterns, making features sensitive to multilevel motions and possibly leading to better performance.
Method $<\delta^{x}_{avg}$ Inference Time
DINOv2 (ViT-S/14) [36] 39.8 0.194
DINOv2 (ViT-B/14) [36] 41.9 0.443
Chrono (ViT-S/14) 68.0 0.575
Chrono (ViT-B/14) 70.1 1.197
Table 7: Inference Time Comparison with DINOv2. We measure inference time for processing 24 frames without point prediction.
Inference time comparison.
Table 7 compares our inference time with that of DINOv2 [36], measuring the time taken to extract features for 24 frames. While our method exhibits slower inference due to temporal injection into the feature, Chrono achieves a significantly higher score than vanilla DINOv2. This higher accuracy compensates for the increased time required for feature extraction.

### v2

@@ Abstract
Figure 1: Chrono is a highly precise, temporally-aware feature backbone specifically designed for point tracking. Unlike prior methods that rely on heavy iterative refinement for each query point [19, 12, 13, 11, 25] or test-time optimization [46, 49], Chrono achieves competitive tracking performance through simple feature matching using a non-parametric operation. Moreover, its performance can be further improved by an optional refinement step.
@@ 1 Introduction
Figure 2: Comparison of initial trajectories from Chrono and existing feature backbone for point tracking. Chrono demonstrates a significant improvement over existing feature backbones such as TSM-ResNet [32].
@@
Despite its critical role, temporally-aware backbone has been relatively under-explored in this area. Early efforts, such as simply incorporating TSM-ResNet [32] in point tracking models [12, 13], yielded limited improvements. This outcome may be attributed to two factors: First, training on small synthetic datasets from scratch [16] may limit generalization to real-world data. Second, TSM-ResNet’s temporal awareness, constrained to adjacent frames, may not provide long-term temporal context for point tracking.
@@
To overcome these limitations, we introduce a temporally-aware feature backbone for point tracking, called Chrono, as illustrated in Figure 1. Our key insights are twofold: First, we leverage robust feature representations learned from large-scale, real-world datasets. We employ DINOv2 [37], a pre-trained model known for strong feature representations across various tasks such as segmentation, classification, and localization [30, 37]. However, these representations are not directly compatible with point tracking due to a lack of temporal awareness. Second, we incorporate temporal awareness by designing a temporal adapter that enables the pre-trained backbone to process data with temporal context, without losing its learned knowledge. This enables the use of these strong feature backbones for point tracking tasks. Unlike the adjacent-frame context used in TSM-ResNet [32], our approach incorporates a temporal context that is six times longer, enabling the capture of complex dynamics and enhancing tracking performance, as shown in Figure 2.
Experimental results show that tracks estimated from Chrono features, computed solely using soft argmax without any learnable layers or additional temporal information after feature extraction, significantly outperform traditional feature backbones commonly used in point tracking as well as DINOv2 [37]. Specifically, Chrono achieves a +20.6%p increase in the position accuracy on the TAP-Vid-DAVIS [12] dataset compared to TSM-ResNet-18 [32]. Additionally, our method demonstrates superior efficiency and comparable performance compared to point tracking models equipped with refiners that inject temporal information post-feature extraction. Despite the absence of learnable layers after feature extraction, Chrono achieves high precision, underscoring that directly embedding temporal information within the features is both efficient and powerful. Optionally, we can add iterative refinement on top of our backbone, which can further enhance the performance of the refiner compared to when an existing backbone is used.
@@ 2 Related Work
With recent advances in large-scale self-supervised training [22, 37], effectively transferring its vast knowledge has become a key challenge. Specifically, the strong features of DINOv2 have been adapted to various tasks. [43] fine-tunes DINO for video tasks in a self-supervised manner. Additionally, DINO has been applied to various correspondence tasks [1, 35, 44, 30, 17] and has demonstrated robustness in establishing correspondence and semantic segmentation [18, 1, 36], showing its rich semantics. We focus on adapting this representation to the point tracking task, enhancing its temporal consistency while preserving its pre-trained knowledge.
@@ 3.1 Preliminaries and Motivation
Point tracking [31, 13, 14, 12, 25, 11, 10] in videos involves establishing correspondences of specific points across frames, determining their positions over time, and identifying whether they are visible or occluded. Formally, given a video sequence $\{\mathcal{I}_{t}\}_{t=0}^{T-1}$, where $\mathcal{I}_{t}\in\mathbb{R}^{H\times W\times 3}$ represents the $t$-th frame of height $H$ and width $W$, and a query point $q=(x_{q},y_{q},t_{q})\in\mathbb{R}^{3}$ at frame $t_{q}$, the goal is to produce a trajectory $\mathcal{T}=\{\hat{p}_{t}\}_{t=0}^{T-1}$, where $\hat{p}_{t}\in\mathbb{R}^{2}$ denotes the estimated position of the point at time $t$, and an associated visibility probability $\mathcal{V}=\{v_{t}\}_{t=0}^{T-1}$, where $v_{t}\in[0,1]$ indicates the likelihood of the point being visible at frame $t$. In this work, we focus on predicting accurate positions rather than occlusions, as our primary goal is to achieve better point correspondences.
Motivation.
@@ 3.2 Temporal Adapter
$\alpha^{(t,n)}(x,y)=\frac{\exp\left(\mathbf{Q}_{t}(x,y)\cdot\mathbf{K}_{t+n}(x,y)/\sqrt{d_{k}}\right)}{\sum_{n^{\prime}=-k}^{k}\exp\left(\mathbf{Q}_{t}(x,y)\cdot\mathbf{K}_{t+n^{\prime}}(x,y)/\sqrt{d_{k}}\right)},$ (3)
where $d_{k}$ is a scaling factor [47], $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are projections of $\mathbf{f}^{\text{down}}$, obtained through the linear projection layers $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$, and $\mathbf{W}_{V}$.
@@ 3.3 Point Prediction with the Feature Backbone
To estimate the position of points in frame $t$, we apply the soft-argmax operation [29] to the correlation map $C_{t}$. The soft-argmax computes a weighted average of all spatial positions, with weights given by the softmax of the correlation values:
@@
$\mathcal{L}_{\text{Huber}}(\hat{{p}}_{t},{p}_{t})=\begin{cases}\frac{1}{2}\left\|\hat{{p}}_{t}-{p}_{t}\right\|^{2},\!&\text{if }\left\|\hat{{p}}_{t}-{p}_{t}\right\|\leq\delta,\\ \delta\cdot\left(\left\|\hat{{p}}_{t}-{p}_{t}\right\|-\frac{1}{2}\delta\right),&\text{otherwise},\end{cases}$ (9)
where ${p}_{t}$ is the ground truth position, and $\delta$ is a threshold parameter. For occluded points indicated by the ground truth visibility status $v_{t}$, we exclude the loss computation, effectively masking out those time steps:
$\mathcal{L}_{t}=(1-v_{t})\cdot\mathcal{L}_{\text{Huber}}(\hat{{p}}_{t},{p}_{t}).$ (10)
@@ 4 Experiments
Backbone RGB-Stacking-Strided Kinetics-Strided DAVIS-Strided
$<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$
ResNet-18 [20, 13] 36.6 64.6 82.7 90.9 94.7 73.9 10.5 35.3 65.7 81.3 88.6 56.3 9.7 31.1 60.6 78.3 87.0 53.3
TSM-ResNet-18 [32, 12] 18.1 55.5 83.7 89.8 92.5 67.9 9.15 33.2 64.6 79.2 86.5 54.5 8.2 26.8 53.6 73.5 83.9 49.2
CoTraker [25] Backbone 36.9 60.1 74.9 82.5 87.6 68.4 31.0 31.0 50.5 70.6 74.9 58.2 33.7 52.5 63.4 68.3 72.0 58.0
DINOv2 (ViT-S/14) [37] 3.2 13.6 44.9 77.4 87.6 45.3 4.9 15.8 41.9 73.8 86.0 44.5 7.3 22.7 52.9 80.0 89.5 50.4
DINOv2 (ViT-B/14) [37] 4.2 15.2 46.6 78.9 88.1 46.6 5.9 18.6 45.9 75.7 86.9 46.6 10.0 28.6 59.5 82.9 87.4 54.4
Chrono (ViT-S/14) 65.5 81.7 88.8 92.3 93.7 84.3 32.2 55.2 73.9 84.0 88.7 66.8 29.7 56.4 76.7 86.8 90.9 68.0
Chrono (ViT-B/14) 68.3 84.0 90.3 92.9 94.5 86.0 33.5 57.2 75.8 85.8 90.2 68.5 31.8 59.2 78.8 88.4 92.2 70.1
Table 1: Quantitative comparison on the TAP-Vid datasets [12] with the strided query mode. Best scores are in bold and second best are underlined for each column.
Backbone RGB-Stacking-First Kinetics-First DAVIS-First
$<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$ $<\delta^{0}$ $<\delta^{1}$ $<\delta^{2}$ $<\delta^{3}$ $<\delta^{4}$ $<\delta^{x}_{avg}$
ResNet-18 [20, 13] 31.4 58.3 77.8 87.7 92.4 69.6 8.6 28.8 56.5 74.2 83.3 50.3 9.0 27.3 54.9 73.7 84.1 49.8
TSM-ResNet-18 [32, 12] 19.2 51.7 78.7 86.8 90.5 65.4 7.8 28.1 55.2 71.4 80.2 48.6 7.3 23.1 46.7 66.6 79.2 44.6
CoTraker [25] Backbone 32.0 54.8 70.2 78.2 84.0 63.9 24.1 41.9 55.1 62.2 67.0 50.1 27.8 46.4 58.6 64.1 68.5 53.1
DINOv2 (ViT-S/14) [37] 3.0 11.8 38.9 72.2 84.1 42.0 4.2 13.6 36.3 65.9 79.2 39.8 6.0 19.9 47.5 73.7 84.5 46.3
DINOv2 (ViT-B/14) [37] 3.6 13.2 41.1 74.1 84.8 43.4 5.1 16.0 40.0 67.9 80.4 41.9 8.9 24.7 53.8 77.0 87.1 50.3
Chrono (ViT-S/14) 58.2 77.2 85.4 89.1 91.3 80.2 24.8 46.2 65.8 77.5 83.1 59.5 24.0 49.2 71.2 82.8 87.9 63.0
Chrono (ViT-B/14) 61.8 79.6 86.9 90.2 92.2 82.1 26.0 48.4 68.2 79.8 85.3 61.6 26.1 52.6 74.5 84.9 90.0 65.6
Table 2: Quantitative comparison on the TAP-Vid datasets [12] with the query-first mode. Best scores are in bold and second best are underlined for each column.
@@ 4.2 Evaluation Protocol
To assess our approach, we employ the TAP-Vid benchmark [12], which comprises both real and synthetic video sets. The real videos come with accurate annotation tracks, while the synthetic videos are paired with perfect ground-truth trajectories. The benchmark consists of three datasets: TAP-Vid-RGB-Stacking, TAP-Vid-Kinetics, and TAP-Vid-DAVIS. TAP-Vid-RGB-Stacking [28] contains synthetic videos of a robot gripper stacking objects and includes 50 videos. The TAP-Vid-Kinetics [27] consists of 1,189 YouTube videos with diverse difficulties, including intense motion blur and abrupt scene transitions, while the TAP-Vid-DAVIS dataset [39] provides 30 videos featuring challenges such as substantial scale shifts of dynamic object.
@@
We assess the accuracy of the predicted tracks using two metrics: position accuracy at various thresholds and average position accuracy $(<\delta^{x}_{\text{avg}})$. Position accuracy is evaluated at five threshold levels: $<\delta^{0}$, $<\delta^{1}$, $<\delta^{2}$, $<\delta^{3}$, and $<\delta^{4}$, corresponding to accuracies within pixel distances of 1, 2, 4, 8, and 16, respectively. Each $<\delta^{x}$ score represents the percentage of visible ground-truth points whose predicted positions fall within the specified threshold. The $<\delta^{x}_{\text{avg}}$ score is the average across all thresholds.
@@ 4.3 Main Results
We evaluate our method against the backbones [20, 32] commonly used in point tracking [11, 12, 13, 25, 26] and DINOv2 [37] in both the strided query (Table 1) and the first query modes (Table 2). Specifically, we compare with ResNet-18 [20] from TAPIR [13], TSM-ResNet-18 [32] from TAP-Net [12], and the backbone from the pre-trained CoTracker model [25], all of which are trained for point tracking. We report position accuracy to demonstrate the effectiveness of temporal information in our model.
Our small model, Chrono (ViT-S/14) with DINOv2 (ViT-S/14), achieves state-of-the-art performance in position accuracy across most thresholds, outperforming other backbones on $<\delta^{x}_{\text{avg}}$. Chrono (ViT-B/14) with DINOv2 (ViT-B/14) delivers even better results. Across RGB-Stacking, Kinetics and DAVIS datasets, in the strided and first query modes, Chrono (ViT-S/14) and Chrono (ViT-B/14) consistently outperform competitors in PCK accuracy at most thresholds. The high accuracy at $<\delta^{0}$ underscores the precision of Chrono features in point tracking, with further gains at higher thresholds. At $<\delta^{4}$, only Chrono backbones achieve over 90% accuracy on the DAVIS dataset in strided mode. Both Chrono (ViT-S/14) and Chrono (ViT-B/14) reach the highest position accuracy among backbones on average, demonstrating Chrono ’s ability to effectively model spatial and temporal information in videos.
@@ 4.4 Analysis and Ablation Study
Comparison to two-stage point tracking models with iterative refiner.
As shown in Table 3, Chrono, although solely a feature backbone without a refiner, achieves point estimation precision comparable to full pipelines with refiners, while demonstrating significantly higher efficiency in terms of throughput. Unlike other models, which require refiners to inject temporal information for each query point, Chrono embeds temporal awareness directly in the feature backbone. This design substantially boosts throughput: for example, Chrono (ViT-B/14) achieves 12.5$\times$ the speed of TAPIR with only a 3.5%p drop in DAVIS accuracy, and even surpasses TAPIR in RGB-Stacking by 11.4%p. Similarly, Chrono (ViT-S/14) delivers 16.4$\times$ higher throughput. These results demonstrate that Chrono provides a simple yet highly effective solution, offering performance comparable to refiner-based models without requiring iterative processing.
Figure 6: Visualization of features. We visualize the features generated by our model and DINOv2 using PCA. The results show that our model demonstrates improved temporal smoothness and finer-grained feature representation relative to DINOv2.
Method RGB-Stacking-S DAVIS-S Throughput # of Refiner Additional
(points/sec) Params. Params.
RAFT [45] 58.6 46.3 23,405.71 4.2M 4.2M
TAP-Net [12] 72.8 53.1 29,535.98 5.5K 5.5K
TAPIR [13] 74.6 73.6 2,097.32 25.9M 25.9M
PIPs [19] 51.0 59.4 46.43 26.0M 26.0M
Chrono (ViT-S/14) 84.3 68.0 34,396.30 0M 16.2M
Chrono (ViT-B/14) 86.0 70.1 26,139.86 0M 26.0M
Table 3: Comparison on point tracking pipelines using refiners. Chrono achieves competitive or superior accuracy while offering significantly higher throughput, despite not using a heavy iterative refiner. TAP-Vid-RGB-Stacking and TAP-Vid-DAVIS metrics report $<\delta^{x}_{avg}$. Throughput is measured on a 24-frame video using a single NVIDIA RTX 3090 GPU.
@@
Analysis on Chrono as a backbone in existing point tracking pipeline.
@@
To model temporal information in our temporal adapter, we tested two approaches for temporal aggregation: convolution and attention. In Table 5, we present an ablation study where only the temporal aggregation layer within Chrono is modified, comparing 1D convolution, 3D convolution, and 1D attention, with all other components kept constant. While 1D convolution and 1D attention aggregate along the temporal axis only, 3D convolution considers both spatial and temporal dimensions. Due to computational constraints, this study is conducted on Chrono (ViT-S/14).
@@
Early transformer blocks often capture local details, while later blocks tend to focus on broader, global patterns and complex relationships [15, 40]. Therefore, it’s plausible that adapters in the initial blocks may learn local motion cues, while those in later blocks might capture overall motion patterns. Placing adapters only in the early or later blocks might emphasize either local or global motion, respectively. While the alternating configuration shows some improvement, it may not be optimal. Integrating temporal adapters between each block could potentially allow for a more balanced capture of both local and global motion patterns, making features sensitive to multilevel motions and leading to better performance.
Method $<\delta^{x}_{avg}$ Inference Time $\downarrow$ GPU Memory $\downarrow$
(sec) (MiB)
DINOv2 (ViT-S/14) [37] 39.8 0.194 1,460
DINOv2 (ViT-B/14) [37] 41.9 0.443 2,816
Chrono (ViT-S/14) 68.0 0.575 1,550
Chrono (ViT-B/14) 70.1 1.197 2,998
Table 7: Computation and latency comparison with DINOv2. We measure inference time and GPU memory usage for processing 24 frames without point prediction.
Computation and latency comparison.
Table 7 compares our inference time and computation with that of DINOv2 [37], measuring the time and GPU memory taken to extract features for 24 frames. While our method exhibits comparable computation and slower inference due to temporal injection into the feature, Chrono achieves a significantly higher score than vanilla DINOv2. This higher accuracy compensates for the increased time required for feature extraction.